# Robust nonlinear state-feedback control of second-order systems

### v1

@@ IV Experimental control results
The proposed control (3) is experimentally evaluated on the electro-magneto-mechanical actuator system with one translational degree of freedom, see Fig. 5.
@@
The step response of the control (3) (denoted further as NLG owing to the nonlinear gains) is compared once with a standard proportional-derivative (PD) control
@@
and once with the sub-optimal nonlinear damping control [14] (denoted further as SOND). The PD control is first designed so that the unperturbed closed-loop control system has two critically damped real poles (denoted further as $\mathrm{PD}_{1}$), i.e., with an optimal value of $\sigma$ that corresponds to the unity damping ratio.
@@
Then, the output control gain $\gamma$ is enhanced (denoted further as $\mathrm{PD}_{2}$), while allowing also for the conjugate complex pole pair with a transient overshoot. The parameters of all four evaluated controllers are listed in Table I. For a fair comparison, the SOND control is assigned with the same gaining factor $\gamma=750$ as the faster PD-control $\mathrm{PD}_{2}$. The feedback gain of the NLG-control $\gamma=25$ is intentionally set an order of magnitude lower than the feedback gains of the PD and SOND controllers. This is for demonstrating that the NLG-control behaves less aggressive during the transient response and, at the same time, achieves a much higher accuracy at steady-state. All experimentally measured control responses, to the same reference step of $y_{\mathrm{ref}}=8$ mm, are shown opposite each other in Fig. 6.
@@
From the response of NLG-control, one can also recognize a nearly uniform behavior of $x_{2}(t)$ during the whole transient phase, cf. Fig. 2 (a). Note that the corresponding inclination of the $x_{1}(t)$ slope depends on the assigned $\gamma$ parameter. It is also worth noting that a residual non-zero, however minor, steady-state error of the NLG-control is due to the sensor’s bias and noise, while the latter is additionally impacting the used $x_{2}(t)$ signal, which is obtained via differentiation and filtering of the measured $x_{1}(t)$.
@@ V Summary
A novel robust nonlinear state-feedback control for second-order systems was introduced. The control law combines the ideas of the state-dependent nonlinear damping introduced in [14] and the sliding mode control theory, which equips the overall feedback law with an output-sign-dependent term able to compensate for bounded perturbations. The resulted control has only one tuning parameter, the overall output feedback gain, whose selection should be based on the upper bound of the perturbation magnitude, and this closed analytic dependence was shown in the paper. We provided a detailed analysis of uniform asymptotic stability property of the control system and finite-time convergence based on the proposed strict Lyapunov function. The control signal appears to be continuous during the transients, apart from the convergence instant, where the state value reaches the origin (similarly to some high order sliding mode control algorithms). Remarkable features of the proposed control is that it: (i) is equivalent (from the structural viewpoint) to a standard linear PD feedback control and, at the same time, (ii) can guarantee for full compensation of the bounded matched disturbances in the second-order systems, while (iii) providing a global finite-time convergence. The proposed control was already evaluated experimentally on an electro-magneto-mechanical actuator system with considerable bounded disturbances.

### v2

@@ IV Experimental control results
The proposed control (3) is experimentally evaluated on the electro-magneto-mechanical actuator system with one translational degree of freedom, see Fig. 5. The system dynamics can be reduced to the model (1).
@@
The step response of the control (3) (denoted further as NLG due to the nonlinear feedback gains) is compared once with a standard proportional-derivative (PD) control and once with the sub-optimal nonlinear damping control [14] (denoted further as SOND).
The benchmarking linear PD control
@@
is designed first so that the unperturbed closed-loop control system has two critically damped real poles, i.e., with an optimal value of $\sigma$ that corresponds to the unity damping ratio. This control is denoted by $\mathrm{PD}_{1}$. Then, its output feedback gain $\gamma$ is enhanced while allowing also for the conjugate complex pole pair with a corresponding transient overshoot. That PD-control realization is denoted by $\mathrm{PD}_{2}$.
The parameters of all four evaluated controllers are listed in Table I.
@@
For a fair comparison, the SOND control is assigned with the same gaining factor $\gamma=750$ as the faster PD-control $\mathrm{PD}_{2}$. The feedback gain of the NLG-control $\gamma=25$ is intentionally set an order of magnitude lower than the feedback gains of the PD and SOND controllers. This is for demonstrating that the NLG-control can behave less aggressive during the transient response and, at the same time, achieve a much higher accuracy at steady-state. All experimentally measured control responses, to the same reference step of $y_{\mathrm{ref}}=8$ mm, are shown opposite each other in Fig. 6.
@@
From the response of the NLG-control, one can also recognize a nearly uniform behavior of $x_{2}(t)$ during the whole transient phase, cf. Fig. 2 (a). Note that the corresponding inclination of the $x_{1}(t)$ slope depends on the assigned $\gamma$ parameter. It is also worth noting that a residual non-zero, however minor, steady-state error of the NLG-control is due to the sensor’s bias and noise, while the latter is additionally impacting the used $x_{2}(t)$ signal, which is obtained via differentiation and filtering of the measured $x_{1}(t)$. The experimentally obtained signal $x_{2}(t)$ (when using the NLG controller) is also shown in Fig. 7 for the sake of clarity.
Figure 7: The velocity state $x_{2}(t)$ of NLG control experiment.
@@ V Summary
A novel robust nonlinear state-feedback control for second-order systems was introduced. The control law combines the ideas of the state-dependent nonlinear damping introduced in [14] and the sliding mode control theory, which equips the overall feedback law with an output-sign-dependent term able to compensate for bounded perturbations. The resulted control has only one tuning parameter, the overall output feedback gain, whose selection should be based on the upper bound of the perturbation magnitude, and this closed analytic dependence was shown in the paper. We provided a detailed analysis of uniform asymptotic stability property of the control system and finite-time convergence based on the proposed strict Lyapunov function. The control signal appears to be continuous during the transients, apart from the convergence instant, where the state value reaches the origin (similarly to some high order sliding mode control algorithms). Remarkable features of the proposed control are that: (i) it is equivalent (from the structural viewpoint) to a standard linear PD feedback control and, at the same time, (ii) it can guarantee (without additional integral action) for full compensation of the bounded matched disturbances in the second-order systems without overshoot, while (iii) providing a global finite-time convergence. The proposed control was already evaluated experimentally on an electro-magneto-mechanical actuator system with considerable bounded disturbances. Analysis of its robustness abilities against process and measurement noise and development of the methods for its discretization can be considered as directions of future research.